# Dynamical systems defined by polynomials with algebraic properties

### v4

@@ 6 $Saut_{P}$ in the case $\deg P=2,3$
(iii) Note that $D\equiv 1~{}(\mbox{mod}~{}4)$. Hence, $D\leq-3$. Then by (7), we should have $c_{1}=0$ or $c_{1}=\pm 1$. If $c_{1}=0$, then $c_{0}=\pm 1$ and hence, $B=\pm I_{2}$. If $c_{1}=\pm 1$, then $D=-4$. Therefore, $Saut_{P}$ is trivial except for the case $D=-4$. If $D=-4$, $Saut_{P}$ contains other elements comming from the case $c=\pm 1$. In fact, in this case, we have $c_{1}=\pm 1,~{}c_{0}=\pm\frac{a_{1}}{2}$ ($\pm$ correspondingly). Note that since $D=a_{1}^{2}-4a_{2}=-4$, $a_{1}$ is even. Hence, $Saut_{P}$ consists of ±$\pm I_{2}$ and ±$\pm\left(\begin{array}[]{cc}\frac{1}{2}a_{1}&1\\ -a_{2}&-\frac{1}{2}a_{1}\end{array}\right)$. Since $\left(\begin{array}[]{cc}\frac{1}{2}a_{1}&1\\ -a_{2}&-\frac{1}{2}a_{1}\end{array}\right)^{2}=I_{2}$, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2. ∎
@@
In the case of $D>0$, there are three sub-cases: i) $D$ non-square; ii) $D$ square, D = 4; iii) $D$ square, $D\neq 4$.
For i), Let $P(z)=z^{2}-3z+1\in\mathbb{Z}[z]$, then $D=5>0$ and non-square. It is clear that $\theta=\frac{\sqrt{5}+3}{2}$ satisfying $\theta^{2}-3\theta+1=0$, and has continued fraction expansion $\theta=[2;1,1,\cdots]$. Then, we have
$\frac{C}{G}=\frac{2}{1},~{}\frac{C^{\prime}}{G^{\prime}}=\frac{1}{0},~{}\frac{E}{F}=\frac{1}{1},~{}\frac{E^{\prime}}{F^{\prime}}=\frac{1}{0}$
and
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}2&1\\ 1&0\end{array}\right)\left(\begin{array}[]{cc}1&1\\ 1&0\end{array}\right)\left(\begin{array}[]{cc}2&1\\ 1&0\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&-1\\ 1&-1\end{array}\right).$
Noting that $a_{2}=-1,~{}a_{1}=3$ and the solution of (9) is $v=w=1$, the solution of (7) is $B=\left(\begin{array}[]{cc}-1&1\\ -1&2\end{array}\right)$ and $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$.
Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\end{array}\right)=\left(\begin{array}[]{cc}-1&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}-x_{0}+x_{1}\\ -x_{0}+2x_{1}\end{array}\right).$
Since $y_{2}=-y_{0}+3y_{1}=-2x_{0}+5x_{1}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{cc}-1&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ -x_{0}+3x_{1}\end{array}\right)=\left(\begin{array}[]{c}-x_{0}+2x_{1}\\ -2x_{0}+5x_{1}\end{array}\right)$
as expected.
For ii), let $P(z)=8z^{2}+6z+1\in\mathbb{Z}[z]$, then $D=4>0.$ $B=\left(\begin{array}[]{cc}3&1\\ -8&-3\end{array}\right).$ Clearly, we have $\det B=-1$ and $B^{2}=I_{2}$. So $Saut_{P}/\{\pm I_{2}\}$ is an finite cyclic group of order 2.
For iii), let $P(z)=-4z^{2}+3z+1\in\mathbb{Z}[z]$, then $D=25$, which is a square and not equal to 4. $B=\pm I_{2}$, so the group $Saut_{P}/\{\pm I_{2}\}$ is trivial.
@@
Let $P(z)=-3z^{2}+1$ with $D=12$. Then $\theta_{1}=\frac{1}{\sqrt{3}}$ is a root of $P(z)=0$ having the continued fraction expansion $\theta_{1}=[0;1,1,2,1,1,2,\cdots]$. Then, we have
$\frac{C}{G}=\frac{0}{1},~{}\frac{C^{\prime}}{G^{\prime}}=\frac{1}{0},~{}\frac{E}{F}=\frac{5}{3},~{}\frac{E^{\prime}}{F^{\prime}}=\frac{1}{0}$
and
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\begin{array}[]{cc}5&1\\ 0&3\end{array}\right)\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&1\\ 3&2\end{array}\right).$
Noting that $a_{2}=3,~{}a_{1}=0$ and $v=1,~{}w=4$ in (9), the solution of (8) is obtained as $B=\left(\begin{array}[]{cc}2&-1\\ -3&2\end{array}\right)$ and $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$.
Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\end{array}\right)=\left(\begin{array}[]{cc}2&-1\\ -3&2\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}2x_{0}-x_{1}\\ -3x_{0}+2x_{1}\end{array}\right).$
Since $y_{2}=3y_{0}=6x_{0}-3x_{1}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{cc}2&-1\\ -3&2\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ 3x_{0}\end{array}\right)=\left(\begin{array}[]{c}-3x_{0}+2x_{1}\\ 6x_{0}-3x_{1}\end{array}\right)$
@@
Let $P(z)=z^{2}-2z+1$ with $D=0$. Then by Theorem 5, $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$ with $B=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)$. Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}x_{1}\\ -x_{0}+2x_{1}\end{array}\right).$
Since $x_{2}=-x_{0}+2x_{1},~{}y_{2}=-y_{0}+2y_{1}=-2x_{0}+3x_{1}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ -x_{0}+2x_{1}\end{array}\right)=\left(\begin{array}[]{c}-x_{0}+2x_{1}\\ -2x_{0}+3x_{1}\end{array}\right)$
as expected.
Example 6.
In the case of $D<0$, there are two sub-cases:
$(i)~{}D=-4\mbox{ and }(ii)~{}D\neq-4.$
For (i), take $P(z)=z^{2}+1\in\mathbb{Z}[z]$, then $D=-4<0.$ It is easy to show that $B=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$ we have $\det B=1$ and $B^{2}=I_{2}$. So $Saut_{P}/\{\pm I_{2}\}$ is an finite cyclic group of order 2.
For (ii), choose $P(z)=z^{2}+3z+3\in\mathbb{Z}[z]$, then $D=-3<0.$ As before, $B=\pm I_{2}.$ So $Saut_{P}/\{\pm I_{2}\}$ is an trivial group.
@@
(i) Assume that the roots $\theta_{1},\theta_{2},\theta_{3}$ of $P(z)=a_{3}z^{3}+a_{2}z^{2}+a_{1}z+1\in\mathbb{Z}[z],a_{3}\neq 0$ are simple, then $B\in Saut_{P}$ if and only if
@@
$\begin{split}~{}~{}~{}\det B&=c_{0}^{3}-a_{3}c_{1}^{3}+a_{3}^{2}c_{2}^{3}-a_{1}c_{0}^{2}c_{1}+(a_{1}^{2}-2a_{2})c_{0}^{2}c_{2}+a_{2}c_{0}c_{1}^{2}\\ &+(a_{2}^{2}-2a_{1}a_{3})c_{0}c_{2}^{2}+a_{1}a_{3}c_{1}^{2}c_{2}-a_{2}a_{3}c_{1}c_{2}^{2}+(3a_{3}-a_{1}a_{2})c_{0}c_{1}c_{2}\\ &=\pm 1.\end{split}$
@@
(ii) From $BM_{P}=M_{P}B$, the expression of B in (i) follows. Or applying (i) with $a_{1}=a_{2}=0,a_{3}=-a.$
@@ 7 Streams over Galois fields
In the last section, We investigate the streams over the Galois field $GF(q^{k})=GF(q)(\theta)$, where $\theta$ is a root of an irreducible polynomial $P(z)\in GF(q)[z]$ of degree $k$. The zeros $\Omega_{P}^{q}$ of $ST(GF(q^{k}))$ by the action of $P(Z)$ in the above sense is discussed. Note that any element in $\Omega_{P}^{q}$ has cycle $q^{k}$, any automorphism of $\Omega_{P}^{q}$ is an strong automorphism, and hence, is represented by a $k\times k$ matrix with entries in $GF(q)$.
@@
Let $P(z)\in GF(q)[z]$ be an irreducible polynomial of degree $k\geq 1$. Let $\Omega_{P}^{q}$ be the zeros of the $P(z)$-action to $St(GF(q))$. Then, the group of the automorphisms of $\Omega_{P}^{q}$ commuting with the shift is a cyclic group of order $q^{k}-1$.
@@
Let $P(z)=a_{0}+a_{1}z+\cdots+a_{k}z^{k}$ with $a_{i}\in GF(q)$. We may assume $a_{0}=1$ since $P(z)$ is irreducible and $GF(q)$ is a field. Then, any element in $\Omega_{P}^{q}$ is cyclic with period $q^{k}-1$, since for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}^{q}$, $x_{n}\in GF(q)~{}(\forall n\in\mathbb{Z})$ is determined by $(x_{n-k},\cdots x_{n-1})\in GF(q)^{k}$. Hence, any automorphism of $\Omega_{P}^{q}$ commuting with the shift can be written as a $k\times k$-matrix with entries in GF(q).
Let $\theta_{1},\cdots,\theta_{k}$ be the roots of $P(z)=0$. Let $\eta$ be a primitive element of $GF(q^{k})=GF(q)(\theta_{1},\cdots,\theta_{k})$. That is, $\eta^{i}=1$ if and only if $i\equiv 0~{}(mod~{}q^{k}-1)$. Then, $\eta$ can be written as $\eta=c_{1}\theta_{1}+\cdots+c_{k}\theta_{1}^{k}$ with $c_{1},\cdots,c_{k}\in GF(q)$ with $c_{1}\cdots c_{k}\neq 0$ Let
@@
Let the matrix $B$ be defined by (5). Then, it is a matrix with entries in $GF(q)$ since the entries are symmetric functions of $\theta_{1},\cdots,\theta_{k}$. Also, it commutes with $M_{P}$. Therefore, it gives an automorphism of $\Omega_{P}^{q}$ commuting with the shift. Moreover, since $\lambda_{1}=\eta$ has the multiplicative order $q^{k}-1$, we have $B^{i}=I_{k}$ if and only if $i\equiv 0~{}(mod~{}q^{k}-1)$. This implies that the set of automorhisms of $\Omega_{P}^{q}$ commuting with the shift is a cyclic group of order $q^{k}-1$, since the cardinality of the set of $B$ as (5) which is nonsingular is $q^{k}-1$. ∎
@@
Let us consider $P(z)=z^{2}+1\in GF(3)[z]$. Let ±$\pm\theta\in GF(9)$ be the roots of $z^{2}+1=0$. Then, ±$\pm\theta+1$ are the primitive elements of $GF(9)$. Then, the following $B$ generates the automorphism group of $\Omega_{P}^{3}$ commuting with the shift.
$B=\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)\left(\begin{array}[]{cc}\theta+1&0\\ 0&-\theta+1\end{array}\right)\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&1\\ 2&2\end{array}\right)$
$012$
@@
The $\sigma$-orbits of $\Omega_{P}^{3}$ and the $B$-orbit of $\Omega_{P}^{3}$ are as above.
Acknowledgements
The authors thank Prof. Shigeki Akiyama and Prof. Hajime Kaneko (Tsukuba Univ., Japan) for their useful suggestions. The simple proof of (iii) of Theorem 7 was taught by Prof. Shigeki Akiyama. The first author was supported by TianYuan Visiting Scholar Program of NSFC(Grant No:12426661).

### v5

@@ 7 $Saut_{P}$ in the case $\deg P=2,3$
(iii) Note that $D\equiv 1~\mbox{or}~0~(\rm{mod}~4)$. Hence, if $D<0$, then $D\leq-3$. Let $B\in Saut_{P}$. By (i-3), $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$ for some $c_{0},c_{1}\in\mathbb{Z}$ with (19). Therefore, $(2c_{0}-a_{1}c_{1})^{2}-Dc_{1}^{2}=4$. This always has a trivial solution that $c_{1}=0,~c_{0}=\pm 1$, corresponding to $B=\pm I_{2}$. We look for the other solutions.
There are 2 cases where the nontrivial sdolutions exist.
(Case 1) D=3D=-3 and $\left\{\begin{array}[]{c}2c_{0}-a_{1}c_{1}=\pm 1\\ c_{1}=\pm 1\end{array}\right.~\mbox{($\pm$ independently)}$.
In this case, $a_{1}$ is an odd number.
If $c_{1}=1$, then $B\in Saut_{P}$ is one of the following:
$B_{1}=\left(\begin{array}[]{cc}\frac{a_{1}+1}{2}&a_{0}\\ -a_{2}&\frac{-a_{1}+1}{2}\end{array}\right)~\mbox{or}~B_{2}=\left(\begin{array}[]{cc}\frac{a_{1}-1}{2}&a_{0}\\ -a_{2}&\frac{-a_{1}-1}{2}\end{array}\right).$
If $c_{1}=-1$, then $B\in Saut_{P}$ is one of the following:
$B_{3}=\left(\begin{array}[]{cc}\frac{-a_{1}-1}{2}&-a_{0}\\ a_{2}&\frac{a_{1}+1}{2}\end{array}\right)~\mbox{or}~B_{4}=\left(\begin{array}[]{cc}\frac{-a_{1}+1}{2}&-a_{0}\\ a_{2}&\frac{a_{1}-1}{2}\end{array}\right).$
It is easy to check using $a_{1}^{2}-4a_{0}a_{2}=-3$ that
$B_{1}^{2}=B_{2},~B_{1}^{3}=-I_{2}~\mbox{and}~B_{3}=-B_{1},~B_{4}=-B_{2}.$
This implies that $Saut_{P}/\{\pm I_{2}\}$ is cyclic group of order 3.
(Case 2) D=4D=-4. In this case, just same as the case $D=4$, we have $Saut_{P}/\pm 1$ is a cyclic group of order 2. ∎
@@
Let $P(z)=5z^{2}-2$ and $\theta=\frac{\sqrt{10}}{5}$ be one of the roots of $P(z)=0$. It has the continued fraction expansion
$\theta=[0;1,\underbrace{1,1,2},\underbrace{1,1,2},\underbrace{1,1,2},\cdots].$
Then let
$\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}1&0\\ 1&1\end{array}\right)$
$\frac{E}{F}:=[1;1,2]=1+\frac{1}{1+\frac{1}{2}}=\frac{5}{3}~,~~\frac{E^{\prime}}{F^{\prime}}:=[1;1]=1+\frac{1}{1}=\frac{2}{1}.$
Then, we have
$Saut_{P}/\{\pm I_{2}\}=\left\{\left(\begin{array}[]{cc}1&0\\ 1&1\end{array}\right)\left(\begin{array}[]{cc}5&2\\ 3&1\end{array}\right)^{n}\left(\begin{array}[]{cc}1&0\\ 1&1\end{array}\right)^{-1};~n\in\mathbb{Z}\right\}.$
If $n=1$ in the above, we have $\left(\begin{array}[]{cc}3&2\\ 5&3\end{array}\right)$. Applying it to
$(\cdots;\frac{3}{7},\frac{5}{7},\frac{4}{7},\frac{2}{7},\frac{3}{7},\frac{5}{7},\cdots)\in\Omega_{P},$
we get
$(\cdots;\frac{5}{7},\frac{2}{7},\frac{2}{7},\frac{5}{7},\frac{5}{7},\frac{2}{7},\cdots)\in\Omega_{P}.$
@@
Let $P(z)=z^{2}-2z+1$ with $D=0$. Then by Theorem 8, $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$ with $B=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)$. Applying it to $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$, we get $(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ with
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\end{array}\right)=\left(\begin{array}[]{c}x_{n+1}\\ -x_{n}+2x_{n+1}\end{array}\right).$
Since $y_{n+1}=-x_{n}+2x_{n+1},~y_{n+2}=-y_{n}+2y_{n+1}=-2x_{n}+3x_{n+1}$, we have
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{n+1}\\ -x_{n}+2x_{n+1}\end{array}\right)=\left(\begin{array}[]{c}-x_{n}+2x_{n+1}\\ -2x_{n}+3x_{n+1}\end{array}\right)$
@@
Let $P(z)=z^{2}+1\in\mathbb{Z}[z]$ with D=4D=-4. Then it is easy to show that $Saut_{P}/\{\pm I_{2}\}=\left\{I_{2},\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)\right\}$. Since $\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)^{2}=I_{2}$, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
@@
(i) Assume that the roots $\theta_{1},\theta_{2},\theta_{3}$ of
$P(z)=a^{\prime}_{3}z^{3}+a^{\prime}_{2}z^{2}+a^{\prime}_{1}z+a^{\prime}_{0}\in\mathbb{Z}[z],~a_{0},a_{3}\neq 0$
are simple. We put $a_{3}=\frac{a^{\prime}_{3}}{a^{\prime}_{0}},~a_{2}=\frac{a^{\prime}_{2}}{a^{\prime}_{0}},~a_{1}=\frac{a^{\prime}_{1}}{a^{\prime}_{0}}$. Then $B\in Saut_{P}$ if and only if
@@
$\begin{split}~~~\det B&=c_{0}^{3}-a_{3}c_{1}^{3}+a_{3}^{2}c_{2}^{3}-a_{1}c_{0}^{2}c_{1}+(a_{1}^{2}-4a_{2})c_{0}^{2}c_{2}+a_{2}c_{0}c_{1}^{2}\\ &+(a_{2}^{2}-2a_{1}a_{3})c_{0}c_{2}^{2}+a_{1}a_{3}c_{1}^{2}c_{2}-a_{2}a_{3}c_{1}c_{2}^{2}+(3a_{3}-a_{1}a_{2})c_{0}c_{1}c_{2}\\ &=\pm 1.\end{split}$
@@
(ii) From $BC_{P}=C_{P}B$, the expression of B in (i) follows. Or applying (i) with $a_{1}=a_{2}=0,a_{3}=-a.$
@@ 8 Streams over Galois fields
In the last section, we investigate the streams over the Galois field $GF(q^{k})=GF(q)(\theta)$, where $\theta$ is a root of an irreducible polynomial $P(z)\in GF(q)[z]$ of degree $k$. The zeros $\Omega_{P}^{q}$ of $ST(GF(q^{k}))$ by the action of $P(Z)$ is discussed. Note that any element in $\Omega_{P}^{q}$ is cyclic with period $q^{k}-1$.
@@
Let $P(z)\in GF(q)[z]$ be an irreducible polynomial of degree $k\geq 1$. Let $\Omega_{P}^{q}$ be the zeros of the $P(z)$-action to $St(GF(q))$. Then, $Saut_{P}$ is a cyclic group of order $q^{k}-1$.
@@
Let
$P(z)=a_{0}+a_{1}z+\cdots+a_{k}z^{k}~\mbox{with}~a_{i}\in GF(q)~(i=0,1,\cdots,k),a_{0},a_{k}\neq 0.$
Then, any element in $\Omega_{P}^{q}$ is cyclic with period $q^{k}-1$, since for any $(x_{n}\in GF(q);n\in\mathbb{Z})\in\Omega_{P}^{q}$, $x_{n}$ is determined by $(x_{n-k},x_{n-k+1},\cdots,x_{n-1})$.
Let $\theta_{1},\theta_{2},\cdots,\theta_{k}$ be the roots of $P(z)=0$. Since $P(z)$ is irreducible, they differ each other. Let $\eta$ be a primitive root of $GF(q^{k})=GF(q)(\theta_{1},\cdots,\theta_{k})$. That is, $\eta^{i}=1$ if and only if $i\equiv 0~({\rm mod}~q^{k}-1)$. Then, $\eta$ can be written as $\eta=c_{1}\theta_{1}+\cdots+c_{k}\theta_{1}^{k}$ with $c_{1},\cdots,c_{k}\in GF(q)$. Let
@@
Let the matrix $B$ be defined by (16). Then, it is a matrix with entries in $GF(q)$ since the entries are symmetric functions of $\theta_{1},\cdots,\theta_{k}$. Also, it commutes with $C_{P}$, and hence, $B\in Saut_{P}$. Moreover, since $\lambda_{1}=\eta$ has the multiplicative order $q^{k}-1$, we have $B^{i}=I_{k}$ if and only if $i\equiv 0~(mod~q^{k}-1)$. This implies that $Saut_{P}$ is a cyclic group of order $q^{k}-1$ since the cardinality of the set of nonzero $B$ as (16) is at most $q^{k}-1$. ∎
@@
Let us consider $P(z)=z^{2}+1\in GF(3)[z]$. Let $\theta\in GF(9)$ be a root of $z^{2}+1=0$. Then, $\theta+1$ is a primitive roots of $GF(9)$. The following $B$ generates $Saut_{P}$.
$B=\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)\left(\begin{array}[]{cc}\theta+1&0\\ 0&-\theta+1\end{array}\right)\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&1\\ 2&2\end{array}\right).$
012012012012
@@
Then $\sigma$-orbits of $\Omega_{P}^{3}$ and the $B$-orbit of $\Omega_{P}^{3}$ are as above.
Acknowledgment: The authors thank Prof. Hajime Kaneko (Tsukuba Univ., Japan) for his useful suggestions. Actually, Theorem 5 is essentially due to him. The first author was supported by TianYuan Visiting Scholar Program of NSFC(Grant No:12426661).